OpenStudy (anonymous):

2w+3squareroot of w=14

OpenStudy (anonymous):

\[2w+3\sqrt{w}=14\]

OpenStudy (anonymous):

solve by substitution

OpenStudy (anonymous):

w= 4 Don't know how to solve it by substitution

OpenStudy (anonymous):

i got 4 also

OpenStudy (anonymous):

what did you use?

OpenStudy (anonymous):

i got w=4 and w = -7/2

OpenStudy (anonymous):

Just looked at the problem & guessed, then checked my math.

OpenStudy (anonymous):

lol wtf, good guess

OpenStudy (anonymous):

well, if you look at it, you know odds are w has to be a number that has an integer square root. And it has to be pretty low. So, 4 was the first number that came to mind, because 9 is the next square, and it's already too high.